OpenStudy (anonymous):

find the vector v sith the given magnitude and the samd direction as u. magnitude= 5 direction=<3,3>

OpenStudy (amistre64):

well, the given vector direction needs to be a unit vector (length 1) so that we can multiply it by 5 to get it to a length of 5

OpenStudy (amistre64):

other than that, we can determine this is a 45 degree vector that has 1-1-sqrt(2) ratio to begin with when sqrt(2) = 5 that means we multiply the sides by 5/sqrt(5)

OpenStudy (amistre64):

err... 5/sqrt(2) that is

OpenStudy (anonymous):

It says the answer is <5sqrt2/2, 5sqrt2/2> I just do not know how to get to the answer exactly

OpenStudy (amistre64):

the key is knowing how to unit a vector ... if we have a vector of a given length; say its 9 inches long. How would you divide this up into pieces that are 1 inch long?

OpenStudy (amistre64):

9/9 = 1 right? so we would divide it by its own length. this works for any length to get it to a unit value.

OpenStudy (amistre64):

how do we determine the length of a given vector?

OpenStudy (anonymous):

with the distance formula? but i dont even know if that is right

OpenStudy (amistre64):

would i steer you wrong?

OpenStudy (anonymous):

haha im sorry my teacher is terrible!

OpenStudy (amistre64):

use the distance formula on your <3,3> and lets see what we get :)

OpenStudy (anonymous):

sqrt 18

OpenStudy (amistre64):

good, or simplified as 3 sqrt(2) when we want to unit a vector, divide itself by its own length \[\frac{1}{3\sqrt{2}}<3,3>=<\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}>\] right?

OpenStudy (anonymous):

im sorry but where did the 1/sqrt2 come in?

OpenStudy (amistre64):

some books hate it when we have sqrts on the bottom, so they like to rewrite them:\[\frac{1}{\sqrt{2}}*\frac{\sqrt{2}}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (amistre64):

\[\sqrt{18}=\sqrt{9*2}=3\sqrt{2}\] use this to divide off the vector; scale i ... \[\frac{1}{3\sqrt{2}}<3,3>=<\frac{\cancel{3}}{\cancel{3}\sqrt{2}},\frac{\cancel{3}}{\cancel{3}\sqrt{2}}>\]

OpenStudy (anonymous):

oh i get it now! thank you so much!!

OpenStudy (amistre64):

youre welcome :)